OpenStudy (vera_ewing):

Which is larger, the radius of the circle of the cylinder or the side length of the square prism that both have a volume of 210 ft^3 and height of 14 ft? Calculate both measurements and compare.

OpenStudy (vera_ewing):

@goformit100 please help? :)

OpenStudy (goformit100):

Ma'am show your working

OpenStudy (vera_ewing):

???

OpenStudy (goformit100):

Show the calculation for the first measurement

OpenStudy (goformit100):

@uri

OpenStudy (vera_ewing):

I don't know how to do the calculations...

OpenStudy (dan815):

\[Volume of Cylinder = \pi * r^2 *h\] \[Volume of SqPrism =b^2*h\],b=base length of the pyramid

OpenStudy (dan815):

Oh the Prism* i mean

OpenStudy (dan815):

not pyramid

OpenStudy (vera_ewing):

The first one is larger right? Because the square is bigger?

OpenStudy (dan815):